3/5-(6/10*1/3)
another question
2/5-(4/10*1/3)
your help will truly be appreciated

Answer by MithrilKnights(8)

Simplify what is inside the parentheses first.
4/10 = 2/5
2/5 * 1/3 = 2 /15
2/5 - 2/15
We have to find the least common multiple of 5 and 15, and it is 15.
So 2/5 = ?/15
? = 6.
So, we get 6/15 - 2/15 = 4/15
And, 4/15 is your answer